OpenStudy (lawlesscrob):

A pool of possible jurors consists of 14 men and 10 women. How many different juries consisting of 5 men and 7 women are possible? a. 2,704,156 c. 840.000331 b. 0.00033069 d. 240,240

OpenStudy (dumbcow):

\[\left(\begin{matrix}14 \\ 5\end{matrix}\right)*\left(\begin{matrix}10 \\ 7\end{matrix}\right)\]

OpenStudy (lawlesscrob):

thanks